Parth (parthkohli):

\[\Large \text{Round-and-round!}\] This is a tutorial about the circle. See the comment below.

Parth (parthkohli):

Back in the early days people noticed something. The circumference of a circle is 3 times the diameter. Then they saw that it's 3.14 times, then the decimals went deeper and deeper. 3.14159... They named that number as 'pi'. Now, we can say that: \[\Large {c = (pi)d}\] We know that diameter is two times radius. So: \[\Large {c = 2(pi)r} \] One common misconception is: \[\Large {(pi) = {22 \over 7}}\] This is not true. 22/7 is the approximate value of pi. While pi is 3.14159, 22/7 is 3.14285. Generally, we take 22/7 as the value of pi because it is very close to pi, but it is not. There is another common misconception. Pi is irrational, but people say it is rational as pi = 22/7, but now you know how they're wrong :) Let's talk about the area of a circle: The area of a circle is \[\Large {(pi)r^{2}}\] We can easily equate the value of the area if the radius is given or vice versa. Let me give you a word problem and a hint. The area of a circle is 36pi. Calculate its radius. \[\Large {36(pi)} = {(pi)r^{2}}\] It's a big hint and easy if you know how to solve equations.

OpenStudy (lgbasallote):

latex advise... \pi = \(\pi\)

Parth (parthkohli):

I was about to write that ;P